A 710- kg car stopped at an intersection is rear-ended by a 1780- kg truck moving with a speed of 16.0 m/s . Part A If the car was in neutral and its brakes were off, so that the collision is approximately elastic, find the final speed of the truck. v truck = m/s Part B Find the final speed of the car. v car v c a r = nothing.
Answer #1

here,

the mass of car , m1 = 710 kg

mass of Truck , m2 = 1780 kg

initial speed of truck , u2 = 16 m/s

let the final speed of car and truck be v1 and v2

using conservation of momentum

m2 * u = m1 * v1 + m2 * v2

1780 * 16 = 710 * v1 + 1780 * v2 .....(1)

using conservation of kinetic energy

0.5 * m2 * u^2 = 0.5 * m1 * v1^2 + 0.5 * m2 * v2^2

1780 * 16^2 = 710 * v1^2 + 1780 * v2^2 .....(2)

from (1) and (2)

v1 = 22.88 m/s

v2 = 6.88 m/s

a)

the final speed of car is 22.88 m/s

b)

the final speed of truck is 6.88 m/s